# Eau De $Q$-Network: Adaptive Distillation of Neural Networks in Deep Reinforcement Learning

### v1

@@ 1 Introduction
In this work, we propose a novel approach to sparse training that gradually prunes the weights of the neural networks at the agent’s learning pace to finish at a sparsity level that is discovered by the algorithm. This behavior is made possible thanks to the combination of two independent methods, namely Distill $Q$-Network and Adaptive $Q$-Network (Vincent et al., 2025b), gathered in a single algorithm coined Eau De $Q$-Network (EauDeQN). In Distill $Q$-Network (DistillQN), the online network is a pruned version of the target network, thereby using the common temporal-difference loss as a distillation loss (Figure 1, left). While DistillQN is also a novel approach to sparse training introduced in this work, it still relies on a hand-designed pruning schedule. This is why we will mainly focus on EauDeQN. Adaptive $Q$-Network (AdaQN) was originally introduced to tune the RL agent’s hyperparameters (Vincent et al. (2025b), Figure 1, right). When combined with DistillQN, the resulting algorithm considers several online networks with equal or higher sparsity levels than the target network. At each target update, the online network with the lowest cumulated loss, represented with a crown in Figure 2, is chosen as the next target network. Therefore, the ability to select between different sparsity levels according to the value of the cumulated loss synchronizes the pruning schedule to the agent’s learning pace. The target network is then copied and pruned to replace the other networks. Crucially, by keeping the online network with the lowest cumulated loss for the next iteration, the algorithm can keep the sparsity level steady if the cumulated loss increases for higher sparsity levels. This results in an algorithm capable of discovering the final sparsity level as opposed to the current methods, which rely on a hand-designed sparsity level.
@@ 5 Experiments
We evaluate our approach on $10$ Atari games (Bellemare et al., 2013) and $6$ MuJoCo environments (Todorov et al., 2012). We apply EauDeQN on $3$ different algorithms corresponding to $3$ RL settings. We use DQN (Mnih et al., 2015) in an online scenario, Conservative $Q$-Learning (CQL, Kumar et al. (2020)) in an offline scenario, and SAC (Haarnoja et al., 2018) in an actor-critic setting. In each RL setting, we compare our approach to its dense counterpart and to PolyPruneQN since it is state-of-the-art among pruning methods (Graesser et al., 2022; Ceron et al., 2024). We focus on obtaining returns comparable to those of the dense approach while reaching high final sparsity levels. For that, we report the Inter-Quantile Mean (IQM, Agarwal et al. (2021)) of the normalized return and the sparsity levels along with $95\%$ bootstrapped confidence intervals over $5$ seeds for the Atari games and $10$ seeds for the MuJoCo environments. We believe that the number of samples used during training is the main limiting factor for pruning algorithms. This is why we report the number of environment interactions as the $x$-axis, except for the offline experiments where we report the number of batch updates. We use the hyperparameters shared by Ceron et al. (2024) for PolyPruneQN as they demonstrate that their method is also effective on the considered RL setting, i.e., $s_{F}=0.95,n=3,t_{\text{start}}=0.2\cdot t_{\text{final}},$ and $t_{\text{end}}=0.8\cdot t_{\text{final}},$ where $t_{\text{final}}$ corresponds to the training length. For EauDeQN, we fix $U_{\text{max}}=3,S_{\text{max}}=0.01,K=5$ and $M=3$ and discuss these values in Section 5.4. The shared hyperparameters are kept fixed across the methods and are reported in Table 4 and 4. We reduced the set of $15$ games selected by Ceron et al. (2024) and Graesser et al. (2022) for their diversity to $10$ games to minimize computational costs. Figure 10 testifies that this set of games conserves a wide variety in the magnitude of the normalized return. Details on experiment settings are shared in Section A. The individual learning curves for each environment are presented in the supplementary material.
@@ 5.1 Online Q-Learning
Figure 5: EauDeDQN’s sparsity schedule (left) differs across the $3$ tested network sizes. Higher final sparsity levels are reached for larger network sizes at the end of the training (right), showcasing EauDeDQN’s adaptivity.
@@
As the representation capacity of the different network architectures is the same (same convolutional layers), one would desire an adaptive pruning algorithm to prune larger networks more, as compared to smaller networks. Figure 5 (left) shows the sparsity schedule obtained by EauDeDQN along with the hard-coded one of PolyPruneDQN. Interestingly, after following a linear curve, the $3$ EauDeDQN’s sparsity schedules split into $3$ different curves to end at a final sparsity level that is environment-dependent (Figure 5, right). Notably, except for the game Pong, larger final sparsity levels are reached for larger networks, as desired. Figure 11 (top) exhibits similar behaviors where higher final sparsity levels are discovered when more environment interactions are available.
@@ 6 Conclusion and Limitations
Limitations Similarly to AdaQN, EauDeQN requires additional time and memory during training. This is a usual drawback of dense-to-sparse approaches (Graesser et al., 2022). Nonetheless, those additional requirements remain reasonable as quantified in Table 2. Another limitation of our work concerns the actor-critic framework. Our algorithm focuses on pruning the critic only. Nonetheless, it is usually the network of the critic that requires a larger amount of parameters (Zhou et al., 2020; Kostrikov et al., 2021; Graesser et al., 2022; Bhatt et al., 2024). Future work could investigate pruning the actor with a simple hand-designed pruning schedule, as done in Xu et al. (2024), while using EauDeQN to prune the critic.
@@ Appendix A Appendix
Figure 13: We evaluate EauDeSAC on $6$ MuJoCo environments for $1$M with the small network. Left: PolyPruneSAC suffers when its final sparsity level is too high as opposed to EauDeSAC, which yields high returns for all considered values of $U_{\text{max}}$. Right: EauDeSAC remains stable across different values of $S_{\text{max}}$ and population size $K$ (see Equation 3), showcasing its robustness w.r.t. hyperparameter changes. The number of subsampled networks $M$ is set to $\left\lceil\frac{K}{2}\right\rceil$. The default hyperparameters of EauDeQN are indicated with a white cross.

### v2

@@ 1 Introduction
In this work, we propose a novel approach to sparse training that gradually prunes the weights of the neural networks at the agent’s learning pace to finish at a sparsity level that is discovered by the algorithm. This behavior is made possible thanks to the combination of two independent methods, namely Distill $Q$-Network and Adaptive $Q$-Network (Vincent et al., 2025b), gathered in a single algorithm coined Eau De $Q$-Network (EauDeQN). In Distill $Q$-Network (DistillQN), the online network is a pruned version of the target network, thereby using the common temporal-difference loss as a distillation loss (Figure 1, left). While DistillQN is also a novel approach to sparse training introduced in this work, it still relies on a hand-designed pruning schedule. This is why we will mainly focus on EauDeQN. Adaptive $Q$-Network (AdaQN) was originally introduced to tune the RL agent’s hyperparameters (Vincent et al. (2025b), Figure 1, right). When combined with DistillQN, the resulting algorithm considers several online networks with equal or higher sparsity levels than the target network. At each target update, the online network with the lowest cumulated loss, represented with a crown in Figure 2, is chosen as the next target network. Therefore, the ability to select between different sparsity levels according to the value of the cumulated loss synchronizes the pruning schedule to the agent’s learning pace. The target network is then copied and pruned to replace the other networks. Crucially, by keeping the online network with the lowest cumulated loss for the next iteration, the algorithm can keep the sparsity level steady if the cumulated loss increases for higher sparsity levels. This results in an algorithm capable of discovering the final sparsity level as opposed to the current methods, which require multiple training iterations to tune the final sparsity level.
@@ 5 Experiments
We evaluate our approach on $10$ Atari games (Bellemare et al., 2013) and $6$ MuJoCo environments (Todorov et al., 2012). We apply EauDeQN on $3$ different algorithms corresponding to $3$ RL settings. We use DQN (Mnih et al., 2015) in an online scenario, Conservative $Q$-Learning (CQL, Kumar et al. (2020)) in an offline scenario, and SAC (Haarnoja et al., 2018) in an actor-critic setting. In each RL setting, we compare our approach to its dense counterpart and to PolyPruneQN since it is state-of-the-art among pruning methods (Graesser et al., 2022; Ceron et al., 2024). We focus on obtaining returns comparable to those of the dense approach while reaching high final sparsity levels. For that, we report the Inter-Quantile Mean (IQM, Agarwal et al. (2021)) of the normalized return and the sparsity levels along with $95\%$ bootstrapped confidence intervals over $5$ seeds for the Atari games and $10$ seeds for the MuJoCo environments. We believe that the number of samples used during training is the main limiting factor for pruning algorithms. This is why we report the number of environment interactions as the $x$-axis, except for the offline experiments where we report the number of batch updates. We use the hyperparameters shared by Ceron et al. (2024) for PolyPruneQN as they demonstrate that their method is also effective on the considered RL setting, i.e., $s_{F}=0.95,n=3,t_{\text{start}}=0.2\cdot t_{\text{final}},$ and $t_{\text{end}}=0.8\cdot t_{\text{final}},$ where $t_{\text{final}}$ corresponds to the training length. For EauDeQN, we fix $U_{\text{max}}=3,S_{\text{max}}=0.01,K=5$ and $M=3$ and discuss these values in Section 5.4. The shared hyperparameters are kept fixed across the methods and are reported in Table 4 and 4. We reduced the set of $15$ games selected by Ceron et al. (2024) and Graesser et al. (2022) for their diversity to $10$ games to minimize computational costs. The subset of $10$ games was selected to maintain a wide variety in the magnitude of the normalized return as shown in Figure 10. Details on experiment settings are shared in Section A. The individual learning curves for each environment are presented in the supplementary material.
@@ 5.1 Online Q-Learning
Figure 5: EauDeDQN’s sparsity schedule (left) differs across the $3$ tested network sizes. Higher final sparsity levels are reached for larger network sizes at the end of the training (right), showcasing EauDeDQN’s adaptivity. The shaded region indicates the variability across seeds.
@@
As the representation capacity of the different network architectures is the same (same convolutional layers), one would desire an adaptive pruning algorithm to prune larger networks more, as compared to smaller networks. Figure 5 (left) shows the sparsity schedule obtained by EauDeDQN along with the hard-coded one of PolyPruneDQN. Interestingly, after following a linear curve, the $3$ EauDeDQN’s sparsity schedules split into $3$ different curves to end at a final sparsity level that is environment-dependent (Figure 5, right). We stress that, similarly to PolyPruneDQN, a baseline following a hard-coded linear schedule would also rely on an accurate tuning of its final sparsity level. Notably, except for the game Pong, larger final sparsity levels are reached for larger networks, as desired. Figure 11 (top) exhibits similar behaviors where higher final sparsity levels are discovered when more environment interactions are available.
@@ 6 Conclusion and Limitations
Limitations EauDeQN requires additional time and memory during training. This is a usual drawback of dense-to-sparse approaches (Graesser et al., 2022). Importantly, Table 2 testifies that training PolyPruneQN with only $2$ different final sparsity levels requires significantly more resources than a single EauDeQN training. Another limitation of our work concerns the actor-critic framework, as it only focuses on pruning the critic. Nonetheless, it is usually the network of the critic that requires a larger amount of parameters (Zhou et al., 2020; Kostrikov et al., 2021; Graesser et al., 2022; Bhatt et al., 2024). Future work could investigate pruning the actor with a simple hand-designed pruning schedule, as done in Xu et al. (2024), while using EauDeQN to prune the critic.
@@ Appendix A Appendix
Figure 13: We evaluate EauDeSAC on $6$ MuJoCo environments for $1$M with the small network. Left: PolyPruneSAC requires tuning as its performance depends on its hard-coded final sparsity level. Conversely, EauDeSAC avoids unsafe final sparsity levels by discovering its final sparsity level, therefore requiring only one training to reach a satisfactory outcome. Right: EauDeSAC remains stable across different values of $S_{\text{max}}$ and population size $K$ (see Equation 3), showcasing its robustness w.r.t. hyperparameter changes. The number of subsampled networks $M$ is set to $\left\lceil\frac{K}{2}\right\rceil$. The default hyperparameters of EauDeQN are indicated with a white cross.
@@ Acknowledgments
Carbon Impact
As recommended by Lannelongue & Inouye (2023), we used GreenAlgorithms (Lannelongue et al., 2021) and ML $CO_{2}$ Impact (Lacoste et al., 2019) to compute the carbon emission related to the production of the electricity used for the computations of our experiments. We only consider the energy used to generate the figures presented in this work and ignore the energy used for preliminary studies. The estimations vary between $1.57$ and $1.82$ tonnes of $\text{CO}_{2}$ equivalent. As a reminder, the Intergovernmental Panel on Climate Change advocates a carbon budget of $2$ tonnes of $\text{CO}_{2}$ equivalent per year per person.